zarkam21 (zarkam21):

Solve the system of equations. Enter your answer as an ordered pair. 6x-y=33 6x+3y=45

OpenStudy (anonymous):

Do you want the point where they intersect?

Mehek (mehek14):

eliminate 6x by subtracting

OpenStudy (anonymous):

try multiplying one of the equations by -1....then add the 2 equatioons, this will cause the 6x's to cancel, then you can find y

Mehek (mehek14):

6x-y=33 - 6x+3y=45 ------------- 0 - 4y = -12

Mehek (mehek14):

so y = 3 now substitute in the equation and solve for x

zarkam21 (zarkam21):

(6,3)

Mehek (mehek14):

correct